### 04 Jan 2024 v1

# link prediction; graph neural network; community detection; network analysis; citation graph; deep learning

Abstract
This study introduces an innovative approach that integrates community detection algorithms with Graph Neural Network (GNN) models to enhance link prediction in scientific literature networks. We specifically focus on the utilization of the Louvain community detection algorithm to uncover latent community structures within these networks, which are then incorporated into GNN architectures to predict potential links. Our methodology demonstrates the importance of understanding community dynamics in complex networks and leverages the strengths of both community detection and GNNs to improve predictive accuracy. Through extensive experiments on bipartite graphs representing scientific collaborations and citations, our approach not only highlights the synergy between community detection and GNNs but also addresses some of the prevalent challenges in link prediction, such as scalability and resolution limits. The results suggest that incorporating community-level information can significantly enhance the performance of GNNs in link prediction tasks. This work contributes to the evolving field of network science by offering a novel perspective on integrating advanced machine learning techniques with traditional network analysis methods to better understand and predict the intricate patterns of scientific collaborations.
keywords:
link prediction; graph neural network; community detection; network analysis; citation graph; deep learning
[…]
1 Introduction
A graph is a special structure of data in computer storage that consists of a collection of vertices and a collection of edges. Link prediction is the process of predicting whether there is a possibility of an edge between vertices that are not directly connected to each other in a specific graph structure, and the corresponding weight of that edge Liben2007LinkPrediction . Traditional link prediction methods include Markov chain and machine learning, and have been used for website browsing prediction and literature citation prediction, etc Sarukkai2000MarkovChains Popescul2003StatisticalRelational . However, the two traditional link prediction methods mentioned above are overly dependent on node attributes in the network when modeling, and node attributes in the actual application scenarios are difficult to obtain and cannot be guaranteed to be reliable, so the traditional methods do not have a satisfactory prediction accuracy. Graph embedding techniques offer significant advantages, particularly in their ability to incorporate complex network structures and node relationships more effectively, leading to improved prediction accuracy in various applications Therefore, link prediction algorithms based on graph embedding have gradually become a hot research topic in recent years.
In this paper, we introduce a novel approach that integrates community detection algorithms with Graph Neural Networks (GNN) to enhance the efficiency and accuracy of link prediction in scientific literature networks. Our method utilizes the Louvain algorithm for community detection, which aids in identifying latent community structures within these networks. These structures are then leveraged by GNN models, specifically designed for this task, to predict potential links with higher precision. This integration not only taps into the community dynamics inherent in scientific collaborations but also harnesses the power of GNNs to interpret and utilize these complex network patterns.
The overall approach, succinctly encapsulated in Figure 1, represents a significant step forward in predictive network analysis, particularly in the context of scientific literature where understanding and predicting collaboration patterns can yield valuable insights.
Figure 1: General idea which integrates community detection algorithms with GNN models.
2 Related Work
2.1 Graph Neural Network
The graph structure in computers belongs to a non-Euclidean structure, which is characterized by irregular arrangement, and it is difficult to define its neighboring nodes for a certain point in the data, and the number of neighboring nodes of different nodes is also different, so it is difficult to be directly applied to deep learning methods such as convolutional neural networks and recurrent data networks. The graph embedding model aims to downscale the graph structure, i.e., mapping a highly dense graph into a low-density graph. Graph embedding algorithms are now widely used in the fields of node classification, link prediction, clustering and visualization Bhagat2011NodeClassification .
Due to the different graph structures and different algorithmic principles applied in the embedding process, graph embedding models are mainly categorized as follows. Cai et al. classified graph embedding models into four categories according to the different structures of the input graphs, i.e., matrix decomposition-based graph embedding, deep learning-based graph embedding, edge reconstruction-based graph embedding, graph kernel-based graph embedding, and generative model-based graph embedding Cai2017GraphEmbeddingSurvey . Goyal et al. classified graph embedding into three categories according to the model utilized in the embedding process, i.e., factorization-based graph embedding model, random walk-based graph embedding model, and deep learning-based graph embedding model Goyal2017GraphEmbedding . Chen et al. classified graph embedding models into seven categories by incorporating the type of the graph along with the model utilized in the embedding process as the basis for the classification, i.e., dimensionality reduction-based graph embedding, random walk-based graph embedding, matrix decomposition-based graph embedding, neural network-based graph embedding, graph embedding for large graphs, graph embedding for hypergraphs, and graph embedding based on attentional mechanisms Chen2020GraphRepresentation . Xie et al. classified the dynamic networks into five categories based on whether they are of snapshot type or continuous time type, and in accordance with their different graph embedding strategies, i.e.: eigenvalue decomposition-based graph embedding, Skim-Gram model-based graph embedding, self-encoding-based graph embedding, and graph convolution-based graph embedding Xie2020DynamicNetwork . Xia et al. classified graph embedding algorithms into four categories based on graph learning techniques, i.e., graph embedding for graph signal processing, graph embedding for matrix decomposition, graph embedding for random walk, and graph embedding for deep learning Xia2021GraphLearning . Based on the existing research ignoring the correlation between static and dynamic graphs, Yuan et al. simultaneously classified static and dynamic graph models into five categories based on the algorithmic principles applied in the embedding process, namely: graph embedding based on matrix decomposition, graph embedding based on random walk, graph embedding based on self-encoder, graph embedding based on graph neural network, and graph embedding based on other methods Yuan2022GraphEmbeddingModel .
The following summarizes the current mainstream graph embedding algorithm models at home and abroad, and briefly introduces their principles. Singular value decomposition (SVD) is a typical algorithm for the decomposition of adjacency matrices, and SVD essentially employs the idea of vector orthogonal transformation to decompose matrices. Further, on the basis of generalized SVD (GSVD), combined with asymmetric transfer, the HOPE algorithm is derived Ou2016AsymmetricTransitivity . Node2vec Grover2016Node2vec , as a classical algorithm for random walk, is inspired by the skip-gram algorithm and DeepWalk algorithm in the language model, and artificially introduces two parameters on the basis of obtaining sequences of neighboring nodes via random walk along the balance the breadth-first traversal and depth-first traversal, so as to take into account the homogeneity and structure of the graph. Due to the strong time-varying nature of the network in real problems, the CTDNE algorithm Nguyen2018DynamicNetwork introduces temporal information on the basis of static graph random walk, i.e., random walk based on the temporal path, which is more suitable for dynamic graph scenarios. Based on the feature decomposition of the Laplace matrix, GCN Kipf2016GraphConvolutional simultaneously deals with the node features as well as the connections between nodes to realize a new node representation. Based on GCN, GraphSAGE Hamilton2017InductiveLearning utilizes known nodes to predict unknown nodes through a fixed-size sampling mechanism, and GAT Velikovic2017GraphAttention introduces an attention mechanism to assign weights to nodes so that the input of complete graph data is not required, which is also more in line with real-world scenarios.
In addition, since many real-world problems are not pairwise node relations, concepts such as hypergraphs and heterogeneous graphs have arisen. In order to better study these graph structures, some scholars have also proposed graph embedding algorithms for hypergraphs, e.g., DHNE Ke2017StructuralDeepEmbedding , HGNN Feng2018HypergraphNN , etc., as well as algorithms for heterogeneous graphs, e.g., HGSL Tang2023Hypergraphs , etc.
2.2 Link Prediction
The following key metrics are mainly used in link prediction algorithms, i.e., Area Under the Curve (AUC) metric, Accuracy, Common Neighbor (CN) metric, Adamic-Adar metric, Resource Allocation metric (RA), Preferred Attachment metric (PA), Random Walk with Restart (RWR) metric and Katz metric. Traditional link prediction algorithms can be divided into two categories: unsupervised and supervised algorithms, where unsupervised algorithms are mainly centered around generating functions, node clustering, likelihood analysis, and other related models, while models under supervised algorithms are more diverse. In recent years, many correlations in graph embedding algorithms have been utilized in link prediction, e.g., DeepWalk algorithm, LINE algorithm. In solving practical problems, since the network structure tends to be more complex and diverse, the main research directions of link prediction models are currently focused on temporal network link prediction, heterogeneous network link prediction and directed graph link prediction.
Links in temporal networks change over time and thus have a more complex and diverse dynamic structure. For link prediction in time-series networks, Rahman et al. transformed the link prediction problem into an optimal coding problem for graph embedding through feature representation extraction, and minimized the reconstruction error through gradient descent, and proposed the Dylink2vec algorithm. This method simultaneously considers the topology of the network and the historical changes of the links, and compares to the traditional link prediction methods based on a single topological structure and DeepWalk based link prediction methods, Dylink2vec experiments yielded higher CN, AA and Katz metrics Rahman2018Dylink2vec . Lekui Zhou et al. proposed the DynamicTriad algorithm in conjunction with representational learning, which can simultaneously preserve the network’s structural and dynamic change information, and through the triadic closure process, the nodes disconnected from the triad are connected to each other, in order to capture the dynamic changes at each moment and performs representation learning Zhou2018DynamicNetwork . Wu et al. established the GETRWR algorithm, which draws on the random walk algorithm in graph embedding, and improves the transfer probability in the RWR metrics by applying the node2vec model, and by introducing a restart factor, i.e., the particles have the probability to return to the initial position in the process of random walk and dynamically calculate the transfer probability of the particles to achieve higher indicators such as AUC and RWR Wu2020LinkPrediction .
Nodes and edges in heterogeneous networks have different data types, and many link predictions based on homogeneous networks are often not directly applicable. For link prediction in heterogeneous networks, Dong et al. proposed the metapath2vec algorithm, which firstly uses a random walk model based on metapaths to construct neighboring nodes in the network, and then performs graph embedding operations on them using the skip-gram algorithm, and finally predicts the neighboring nodes of the nodes through a method based on negative sampling of heterogeneous networks. This algorithm can maximize the preservation of the structure and semantics of the heterogeneous network Dong2017Metapath2vec . Fu et al. on the other hand, through the construction of the neural network model used for heterogeneous network representation learning, and then use the Hin2vec algorithm to capture the different node relationships and network structure in it, combined with the meta-path to the nodes, and then the vector feature learning to generate the meta-paths for comparative experiments. This algorithm better preserves the contextual information, including the correlation and differentiation between nodes, compared to traditional heterogeneous network link prediction methods Fu2017HIN2Vec . Jiang et al. improved the graph convolution neural network, which firstly replaces the traditional graph convolution’s transfer layer by a HeGCN layer to make it more suitable for dealing with heterogeneous networks and combines with an adversarial learning layer to obtain node characterization, and then the Adam model is is trained to reduce the value of the loss function, which improves the AUC metrics and PA metrics in link prediction compared to the deep walk and Hin2vec algorithms Jiang2022LinkPrediction .
Edges in directed networks carry directional information, and the directionality of edges is very common in real-world problems, and ignoring the directional information in them will reduce the model predictability. For link prediction in directed networks, Lichtenwalter et al. proposed the PropFlow algorithm based on random walk algorithm, which assigns parameters to potential connections through a baseline, and the objects will choose routes based on the weights when walking while restricting the direction of the object’s walking to the direction of the edges, and the objects do not need to restart or convergence, so the computation time can be reduced Lichtenwalter2010LinkPrediction . Ramesh Sarukkai et al. used Markov chain algorithm in machine learning for link prediction and path research, which can be modeled dynamically according to the historical change condition and combined with eigenvector decomposition to obtain link tours, and this algorithm provides an effective tool for the research of Web path analysis and link prediction Sarukkai2000MarkovChains . Pan et al. applied neural network to the link prediction of directed graph proposed NNLP algorithm, more suitable and combined with the standard particle swarm algorithm to optimize it, and use error back propagation algorithm and BP iterative learning algorithm to train neural network, the training process fusion using CN, AA, PA, Katz and other indicators, thus improving the model’s complementarity and accuracy Pan2018LinkPredictionNN .
3 Datasets
3.1 Data Collection
In this study, we utilized data extracted from the Web of Science database, focusing on literature pertaining to zinc batteries. Our dataset encompasses a comprehensive collection of 10,187 research articles, identified up to November 2023. The majority of these publications, approximately 9,315 papers, were released post-2010, representing over 90 % of the total corpus. Notably, the articles predominantly appear in leading journals within the field, including the Journal of Power Sources, Chemical Engineering Journal, ACS Applied Materials & Interfaces, Journal of the Electrochemical Society, and Journal of Materials Chemistry A, among others.
For the dataset compilation, a specific search query was employed within the Web of Science database. The query was structured as follows: TI=(Zinc) AND TI=(Anode) OR TI=(Zn) AND TI=(Anode) OR TI=(Zn) AND TI=(Batter*) OR TI=(Zinc) AND TI=(Batter*). This search strategy was designed to capture articles where the title (TI) contained the keywords ’Zinc’ or its abbreviation ’Zn’ in conjunction with ’Anode’ or variations of ’Battery’ (indicated by the wildcard ’*’).
3.2 Train-Test Split
A network graph is created from an edge dataset to facilitate a balanced approach to link prediction. The process begins with the construction of a graph using edge data, symbolizing connections between nodes. Subsequently, the dataset is partitioned into training and test subsets.
To address the prevalent issue of class imbalance in network data, a novel step is introduced where non-existent links are generated. This generation continues until their count matches the number of existing links in the training set, thereby ensuring an equitable representation of both link types. The final training set, enriched with these synthetic non-links, and a comparably prepared test set, serve as the foundation for further model training and evaluation. This methodology emphasizes accurate prediction of both the presence and absence of links in a network, a critical aspect in network analysis. The resulting sizes of the oversampled training set and the test set are 296,574 and 74,144, respectively, demonstrating the significant expansion of the dataset for balanced training.
4 Methods
4.1 Heuristic-Based Methods
4.1.1 Overview
Table 1 provides a summarized comparison of various heuristic-based methods for link prediction in terms of their performance metrics: Precision, Recall, F1 Score, and AUC Score. The table lists five methods, from a baseline ’Random Prediction’ to more sophisticated approaches like ’Resource Allocation’. Each method’s efficacy is quantified, with ’Adamic/Adar’ showing the highest precision and ’Common Neighbor’ exhibiting the best balance between precision and recall as reflected by its F1 Score.
Table 1: Heuristic-Based Methods Overview
Method Precision Recall F1 Score AUC Score
Random Prediction
0.5
0.5
0.5
0.5
Common Neighbor
0.844
0.869
0.856
0.854
Jaccard Coefficient
0.854
0.786
0.819
0.826
Adamic/Adar
0.942
0.651
0.77
0.806
Resource Allocation
0.914
0.766
0.834
0.847
4.1.2 Random Prediction
A random prediction method is employed as a baseline model for evaluating the test set. This method involves randomly assigning binary labels (0 or 1) to each element of the test set, effectively simulating a naive approach where predictions are made without any data-driven learning or inference. This technique establishes a foundational benchmark for assessing the efficacy of more sophisticated models.
The outcomes of this random prediction strategy are quantified using standard performance metrics. The results demonstrate equal values across all metrics with a precision, recall, F1 score, and AUC (Area Under the Curve) score, all at approximately 0.5002. These metrics indicate that the model performs at a level akin to random guessing, as expected from this method. The consistency across all metrics underscores the non-discriminatory nature of the random prediction approach. These results set a baseline that more advanced, data-informed models must exceed to be considered valuable in predictive analytics within this domain.
4.1.3 Common Neighbors
In this segment of the analysis, the Common Neighbors Method is implemented as a technique for link prediction. This method operates on the principle that two nodes in a network are more likely to form a link if they share a significant number of common neighbors. The approach involves calculating the number of common neighbors for each pair of nodes in the test set, and a prediction of a link (label 1) is made if the common neighbors count is one or more; otherwise, no link (label 0) is predicted.
The performance of the Common Neighbors Method is evaluated through several metrics, demonstrating notable effectiveness in predicting links. The precision, which measures the correctness of predicted links, is 0.8437, suggesting that a significant proportion of links predicted by the model are actual links. The recall of 0.8691 implies that the method is effective in identifying actual links within the test set. The F1 score, a balance between precision and recall, stands at 0.8562, further attesting to the method’s robustness. Finally, the AUC score of 0.8540 indicates a strong ability to differentiate between the presence and absence of links. These results suggest that the Common Neighbors Method is a reliable tool in network analysis for link prediction, substantially surpassing the baseline established by the random prediction approach.
4.1.4 Jaccard Coefficient
The Jaccard Coefficient Method Jaccard1912 , another approach for link prediction, is based on the similarity between the sets of neighbors of two nodes. The Jaccard coefficient is calculated as the ratio of the size of the intersection to the size of the union of the neighbor sets of two nodes. Predictions are made based on a set threshold; in this case, a link is predicted if the Jaccard score is equal to or exceeds 0.01.
The performance metrics for the Jaccard Coefficient Method show its effectiveness in link prediction. The precision is 0.8537, suggesting that a significant portion of the predicted links are true links. The recall of 0.7862 demonstrates the method’s capability in identifying a majority of actual links. The F1 score, at 0.8186, reflects a balanced measure of precision and recall. Additionally, an AUC score of 0.8257 highlights the method’s proficiency in distinguishing between link presence and absence. These results showcase the Jaccard Coefficient Method as a viable and efficient tool for predicting links in network analysis.
4.1.5 Adamic/Adar Index
The Adamic/Adar Index Method Adamic2003 is implemented for link prediction, focusing on the shared neighbors of two nodes. The method computes the Adamic/Adar index, which accounts for the number of common neighbors and inversely weights these neighbors by the logarithm of their degrees. A threshold is set (0.5 in this case) to decide whether a link is predicted based on the index value.
The method’s performance is summarized as follows: it achieves a precision of 0.9421, a recall of 0.6514, an F1 score of 0.7702, and an AUC score of 0.8057. These metrics indicate that while the method is highly precise in its predictions, its recall is comparatively lower, suggesting a more conservative prediction of links. The overall results demonstrate the method’s efficiency in discerning link presence in the network.
4.1.6 Resource Allocation Index
The Resource Allocation Index Method Zhou2009 , applied here for link prediction, is based on the idea of allocating resources through common neighbors between pairs of nodes. The Resource Allocation (RA) index is calculated for each pair in the test set, taking into account the degree of common neighbors. A predefined threshold (0.01 in this case) is used to determine if a link is predicted based on the RA index score.
The performance metrics for the Resource Allocation Index Method are as follows: a precision of 0.9144, indicating a high degree of accuracy in the predicted links; a recall of 0.7662, showing the method’s ability to identify a substantial portion of actual links; and an F1 score of 0.8337, reflecting a balanced trade-off between precision and recall. The AUC score is 0.8472, suggesting a strong discriminatory ability between link presence and absence.
4.2 Machine Learning Methods
4.2.1 Overview
Table 2 delineates the performance of three machine learning algorithms—Logistic Regression, Random Forest, and XGBoost—on link prediction tasks, evaluated by Precision, Recall, F1 Score, and AUC Score. It highlights that while XGBoost achieves the highest precision, Random Forest outperforms the others in terms of the F1 Score, suggesting a more balanced performance in precision and recall.
Table 2: Machine Learning Methods Overview
Method Precision Recall F1 Score AUC Score
Logistic Regression
0.78
0.709
0.743
0.754
Random Forest
0.838
0.793
0.815
0.82
XGBoost
0.864
0.635
0.794
0.809
4.2.2 Feature Engineering
In the feature engineering stage of this study, a rigorous process is employed to prepare node attributes for network analysis. This involves filling missing values in textual data, one-hot encoding categorical variables, and applying TF-IDF vectorization to text-based attributes. The integration of these methods results in a multi-dimensional feature space represented as a sparse matrix, encompassing 10,187 samples with 2,026 distinct features. This comprehensive dataset, achieved through meticulous preparation and transformation of node characteristics, is instrumental in constructing a robust foundation for the subsequent predictive modeling in network analysis.
4.2.3 Logistic Regression
In this study, a Logistic Regression model Berkson1944 is employed as a machine learning method for link prediction in the network. Logistic Regression, a widely used statistical model for binary classification, is trained on the engineered feature set, which encapsulates the characteristics of node pairs in the network. The model is optimized to predict the likelihood of link existence between nodes based on the features derived from their attributes.
The performance of the Logistic Regression model is evaluated using standard metrics on both training and test datasets. In the training phase, the model achieves a precision of 0.7242, a recall of 0.7220, an F1 score of 0.7231, and an AUC score of 0.7235. For the test set, the model demonstrates a precision of 0.7799, a recall of 0.7090, an F1 score of 0.7428, and an AUC score of 0.7544. These metrics indicate the model’s effectiveness in accurately predicting links, with a notable performance improvement in the test set compared to the training set. This suggests that the Logistic Regression model, trained on the comprehensive feature set, is a robust tool for link prediction in network analysis.
4.2.4 Random Forest
In this phase of the study, a RandomForestClassifier, a powerful ensemble learning method, is utilized for link prediction in the network. The RandomForest model Ho1995 Ho1998 , known for its robustness and ability to handle complex datasets, is trained on the prepared feature set, which includes a diverse range of node attributes and their interactions.
The RandomForest model’s performance is assessed using a variety of metrics. During training, the model demonstrates exceptionally high precision, recall, F1 score, and AUC score, all approximately 0.9986, indicating an almost perfect fit to the training data. However, when evaluated on the test set, the model exhibits a precision of 0.8379, a recall of 0.7926, an F1 score of 0.8146, and an AUC score of 0.8196. These results reflect the model’s effectiveness in generalizing to new data, with a notable distinction between its training and testing performance. The high precision on the test set suggests the model’s strong capability in accurately identifying true links, while the recall indicates its efficiency in capturing a substantial proportion of actual links. These metrics collectively demonstrate the RandomForest model’s efficacy as a robust tool for predicting links in network analysis.
4.2.5 XGBoost
The study employs the XGBoost classifier, an advanced gradient boosting framework, for the task of link prediction in networks. XGBoost ChenGuestrin2016 is renowned for its high efficiency and effectiveness, particularly in dealing with complex datasets. It operates by sequentially building trees, where each new tree attempts to correct the errors made by the previous ones, resulting in a strong predictive model.
In terms of performance, the XGBoost model demonstrates a balanced and effective capability in both the training and testing phases. The training metrics reveal an accuracy of 0.7948, precision of 0.8186, recall of 0.7574, F1 score of 0.7869, and an AUC score of 0.7948. The test metrics exhibit an accuracy of 0.8095, precision of 0.8637, recall of 0.7349, F1 score of 0.7941, and an AUC score of 0.8095. These results indicate the model’s strong performance in accurately predicting links, with particularly high precision in the test set, suggesting its ability to generalize well to new data. The XGBoost model thus proves to be a robust and reliable tool in the context of network analysis for link prediction.
4.3 Graph Neural Networks
4.3.1 Overview
Table 3 shows the comparison between Graph Neural Network (GNN) models with and without the integration of Louvain community detection reveals the impact of incorporating structural community information into link prediction tasks. The results indicate a consistent improvement in model performance when community detection is applied.
For the GAT model, the addition of Louvain community detection leads to increases in precision, recall, F1 score, and AUC score, suggesting that community context enhances the model’s ability to predict links accurately. Similar trends are observed in the GATv2 model, where integrating community features results in a rise in all evaluated metrics, notably in the AUC score.
The GCN model, already exhibiting strong performance, further benefits from the inclusion of community detection, as evidenced by the increase in precision, recall, F1 score, and AUC score. This improvement is also mirrored in the GCNv2 model, although the precision slightly decreases while the other metrics show enhancements.
Lastly, the GraphSAGE model shows considerable improvements across all metrics with the inclusion of Louvain community detection. This suggests that adding community features is particularly effective for models like GraphSAGE, which rely heavily on neighborhood aggregation.
Table 3: Graph Neural Networks Overview
Method Precision Recall F1 Score AUC Score
GAT
0.573
0.918
0.705
0.777
GAT + Louvain
0.606
0.923
0.732
0.823
GATv2
0.582
0.932
0.717
0.812
GATv2 + Louvain
0.608
0.913
0.730
0.830
GCN
0.664
0.957
0.784
0.892
GCN + Louvain
0.682
0.965
0.799
0.906
GCNv2
0.700
0.976
0.815
0.917
GCNv2 + Louvain
0.680
0.970
0.799
0.919
GraphSAGE
0.656
0.939
0.772
0.859
GraphSAGE + Louvain
0.688
0.952
0.799
0.878
4.3.2 Graph Attention Networks
In this study, a Graph Attention Network (GAT) is employed for link prediction, leveraging the power of attention mechanisms in graph neural networks. The GAT model uses multiple attention heads to focus on different parts of the node’s neighborhood, allowing for a more nuanced aggregation of neighbor information. This structure is particularly effective in capturing the complex dependencies within the graph data.
The GAT model achieves a precision of 0.6065 and a recall of 0.9231 in link prediction, with an F1 score of 0.7320 and an AUC score of 0.8225. These metrics underscore its effectiveness in identifying relevant links and its strong overall predictive performance.
The study employs GATv2 Brody2022 , an evolution of the Graph Attention Network (GAT), which incorporates enhanced attention mechanisms for link prediction. GATv2 is characterized by its attention heads that enable more nuanced feature aggregation from a node’s neighborhood, effectively capturing complex interactions within the graph.
In testing, the GATv2 model demonstrates a precision of 0.5824 and a recall of 0.9317, culminating in an F1 score of 0.7168. The AUC score stands at 0.8118, indicating its proficient predictive ability in distinguishing between the presence and absence of links.
Figure 2 includes the training loss data for the GAT and GATv2 models, spanning epochs 0 to 300. The GAT losses are marked with triangles and the GATv2 losses with crosses, using colors consistent with the previous plots.
Figure 2: Training loss over epochs for GAT models.
4.3.3 Graph Convolutional Networks
In this segment of the research, a Graph Convolutional Network (GCN) is implemented for link prediction in network analysis. This GCN includes additional layers and dropout for regularization, enhancing its capability to capture complex patterns in graph data. The model, by operating directly on the graph structure, efficiently aggregates and transforms node features, leveraging the inherent topological information.
The GCN model demonstrates commendable performance in link prediction, achieving a precision of 0.6638, a recall of 0.9570, and an F1 score of 0.7839. Its AUC score of 0.8921 highlights its effective discrimination between link presence and absence, underscoring the model’s robustness in network analysis.
The GCNv2 model Tang2019 , an advanced version of the Graph Convolutional Network (GCN), is implemented for link prediction. The GCNv2 model integrates additional convolution layers and a linear transformation layer, enhancing its capacity to process complex graph structures. Key parameters such as alpha and theta are adjusted to optimize the model’s performance.
The GCNv2 model exhibits impressive results in link prediction, achieving a precision of 0.6796, a recall of 0.9697, and an F1 score of 0.7991. Its AUC score of 0.9199 further demonstrates its strong capability in differentiating between the presence and absence of links in the network.
Figure 3 illustrates the training loss trajectories of two graph neural network models, GCN and GCNv2, over a series of 300 epochs. The GCN model exhibits a gradual decrease in loss with minor fluctuations, while the GCNv2 model shows a more pronounced decrease, particularly in the early epochs.
Figure 3: Training loss over epochs for GCN models.
4.3.4 GraphSAGE
The study implements GraphSAGE, a Graph Neural Network model known for its effectiveness in learning from large graphs. GraphSAGE utilizes neighborhood sampling and aggregating functions to generate node embeddings, efficiently capturing local graph structures.
The GraphSAGE model demonstrates a precision of 0.6556 and a recall of 0.9392 in link prediction, resulting in an F1 score of 0.7722. Its AUC score of 0.8589 indicates a strong capability in distinguishing between the presence and absence of links.
Figure 4 includes the updated training loss data for the GraphSAGE model, spanning epochs 0 to 300.
Figure 4: Training loss over epochs for GraphSAGE model.
4.3.5 Graph Neural Networks with Community Detection
In this phase of the study, community detection is integrated into the preprocessing of graph data for enhanced network analysis. The Louvain method, a popular community detection algorithm, is applied to the network graph to identify communities or clusters within it. This method is advantageous for uncovering the modular structure of the network, revealing groups of nodes that are more densely connected internally than with the rest of the network.
After detecting communities, each node is assigned a community label, which is then transformed into a feature using one-hot encoding. These community features are combined with the existing node features, previously processed through methods like TF-IDF vectorization and one-hot encoding for textual and categorical data, respectively. The resulting combined feature set is more comprehensive, now including both individual node characteristics and their community affiliations.
The enriched dataset, now encapsulating both intrinsic node characteristics and their community contexts, provides a robust foundation for GNN models to analyze network structure and predict links more effectively. This advanced preprocessing approach, by integrating community detection, enhances the model’s capacity to understand and leverage the complex, multi-faceted nature of network data.
Figure  5 offers a comprehensive comparison of training loss trajectories across five graph neural network models—GCN, GCNv2, GAT, GATv2, and GraphSAGE, each enhanced with the Louvain method for community detection, over 300 epochs. The initial loss values vary across models, reflecting different starting efficiencies. Notably, GATv2 exhibits significant early fluctuations, peaking at epoch 10. Despite these variations, all models demonstrate a general trend of decreasing loss, indicating progressive learning and improvement. The use of a sequential color palette and distinct markers (circles for GCN, squares for GCNv2, triangles for GAT, crosses for GATv2, and diamonds for GraphSAGE) aids in differentiating each model’s performance. As epochs advance, the models’ loss values begin to converge, suggesting stabilization in learning. This visualization serves as an insightful tool for analyzing the learning behavior and efficiency of different graph neural network models over an extended training period.
Figure 5: Training loss over epochs for different models.
5 Discussion
5.1 Conclusion
In summary, the integration of Louvain community detection into GNN models consistently enhances their performance in link prediction tasks. The improvements across various metrics underline the significance of incorporating community structure into the analysis, providing a more holistic understanding of the network’s underlying patterns. This approach exemplifies the potential of combining node-specific features with broader network structures to enhance the predictive capabilities of GNN models.
5.2 Limitation
The Louvain community detection algorithm, when applied in Graph Neural Networks for link prediction, may encounter a resolution limit issue. This limitation arises because the algorithm is prone to merging smaller communities into larger ones, thereby potentially overlooking finer, yet significant, community structures. This oversight can lead to suboptimal performance in link prediction tasks where the identification of smaller, more nuanced communities is crucial for accurate predictions.
5.3 Future Work
Future research efforts should focus on further refining and advancing the integration of community detection methods with GNN architectures for link prediction. This includes exploring alternative community detection algorithms that may address the limitations of current methods, such as the resolution limit issue in the Louvain algorithm.
Additionally, enhancing the GNN models to better capture and utilize the nuanced structural information provided by community labels could lead to more accurate and efficient link prediction in complex networks. The exploration of dynamic community detection methods to accommodate evolving network structures, as well as the application of these techniques to a wider range of datasets, would also be beneficial.
[…]
Conceptualization, C.L. and Y.H.; methodology, Y.H and Y.S.; software, Y.H; validation, Y.H.; formal analysis, Y.H.; investigation, Y.H. and C.L.; data curation, C.L. and K.W.; writing—original draft preparation, Y.H. and C.L.; writing—review and editing, Y.H. and C.L.; visualization, Y.H.; supervision, C.L. and S.Y.; project administration, H.X.; All authors have read and agreed to the published version of the manuscript.
[…]
This research was funded by the National Natural Science Funding of China grant number (No.72274113), in part by the Chinese Academy of Science’s “Light of West China” program and the Taishan Scholar Foundation of Shandong province of China (tsqn202103069).
[…]
Not applicable.
[…]
Not applicable.
[…]
The data presented in this study are available on request from the corresponding author.
[…]
The authors declare no conflict of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
XGBoost Extreme gradient boosting
GNN Graph neural network
GAT Graph attention network
GCN Graph convolutional network
GraphSAGE Graph sample and aggregation
[…]